A right cylinder is formed from a rectangular sheet. Its volume is
36/π cubic units. What is the rectangular sheet’s area?
The rectangle's length is 4.

Answer by mananth(16946)

A right cylinder is formed from a rectangular sheet. Its volume is
36/π cubic units. What is the rectangular sheet’s area?
The rectangle's length is 4.

The volume of a right cylinder V = pi*r^2*h where r is the radius and h height
The length of the rectangle is given as 4.
That means the width forms the cicumference of the cylinder when the sheet is rolled.
Circumference = 2*pi*r
2 * pi* r = 4
r= 4/(2*pi)
r= 2/pi
V = pi*r^2*h
insert the value of r in the above equation
V= pi*(2/pi)*(2/pi)* h
36/pi = 4/(pi)^2 * h Given that V= 36/pi
36*(pi)^2 / 4pi= h
9pi = h --h becomes the length of the sheet, L
Area of sheet = L * W
= 9pi * 4
= 36pi square units

M Ananth------------ mananth@hotmail.com